Calculate the electric field inside an insulating sphere of radius R at a distance r from the center


1
Expert's answer
2021-05-10T08:28:08-0400

The electric field inside an insulating sphere of radius R at a distance r from the center is


"E=\frac{1}{4\pi\epsilon_0}·\frac{Qr}{R^3}"